# DF-Calib: Targetless LiDAR-Camera Calibration via Depth Flow

### v1

@@ IV-D Ablation Study
We conduct ablation experiments on the KITTI Odometry and KITTI Raw datasets under an initial perturbation of $(10^{\circ},0.25\mathrm{m})$ to evaluate the effectiveness of our proposed components, using mean absolute translation error and mean absolute rotation error as evaluation metrics. The results are shown in Table III.
@@
Without completing the sparse LiDAR depth map and directly using it with the image-estimated dense depth map $D_{I}^{d}$ via a shared feature encoder, the error approximately tripled, demonstrating that depth completion reduces the information density gap and enhances feature consistency. When processing $D_{L}^{d}$ and $D_{I}^{d}$ with separate encoders instead of a shared one, the error increased by about $10\%$, and model parameters grew from 19M to 28M, highlighting the shared encoder’s efficiency in feature extraction and network complexity reduction. Removing both depth completion and the shared encoder led to an approximate $25\%$ increase in error and a $47\%$ rise in model parameters, reinforcing their importance.
@@
Removing the constraint of the Reliability Map on dense depth features resulted in an error increase of around $15\%$ on both datasets, indicating its role in suppressing noise introduced by depth completion. Replacing the PWSF loss with standard $\mathcal{L}_{1}$ loss approximately doubles the error, demonstrating that PWSF loss effectively filters out invalid points and prioritizes salient regions, improving depth flow estimation and overall accuracy.
V Conclusion
In this paper, we propose DF-Calib, a targetless LiDAR-camera calibration method that reformulates the task as an intra-modality depth flow estimation problem. We unify LiDAR and camera data into a dense depth representation to bridge the modality gap. This facilitates more effective feature matching via a shared feature encoder and utilizes a reliability map to prioritize valid points, thereby enhancing calibration accuracy. Additionally, we introduce a perceptually-weighted sparse flow loss, which improves calibration precision by emphasizing salient regions. Experimental results on multiple datasets demonstrate the effectiveness of DF-Calib, achieving accurate extrinsic calibration and robust generalization across diverse setups. In the future, we will explore the use of depth flow for calibrating other heterogeneous sensors to adapt to diverse robotic platforms.

### v2

@@ 4.2.3 Ablation Study
We conduct ablation experiments on the KITTI Odometry and KITTI Raw datasets under an initial perturbation of $(10^{\circ},0.25\mathrm{m})$ to evaluate the effectiveness of our proposed components, using mean absolute translation error and mean absolute rotation error as evaluation metrics. The results are shown in Tab. 3.
@@
Without completing the sparse LiDAR depth map and directly using it with the image-estimated dense depth map $D_{I}^{d}$ via a unified feature encoder, the error approximately tripled, demonstrating that depth completion reduces the information density gap and enhances feature consistency. When processing $D_{L}^{d}$ and $D_{I}^{d}$ with separate encoders instead of a unified one, the error increased by about $10\%$, and model parameters grew from 19M to 28M, highlighting the unified encoder’s efficiency in feature extraction and network complexity reduction. Removing both depth completion and the unified encoder led to an approximate $25\%$ increase in error and a $47\%$ rise in model parameters, reinforcing their importance.
@@
Removing the constraint of the Reliability Map on dense depth features resulted in an error increase of around $15\%$ on both datasets, indicating its role in suppressing noise introduced by depth completion. Replacing the PWSF loss with a standard $\mathcal{L}_{1}$ loss increases the error by $23\%$, while using an NLL loss leads to a $19\%$ degradation. These results demonstrate that our PWSF loss provides more effective uncertainty-aware supervision by better capturing spatial reliability and suppressing unreliable regions.
Table 4: Runtime Analysis
Depth Generation Network Inference PnP Solver Total
Time [ms] 37 29 25 91($\sim$11Hz)
4.2.4 Runtime Analysis
To evaluate the computational efficiency of our method, we perform a comprehensive runtime analysis, with the result presented in Tab. 4. The initial depth generation stage consists of two concurrent modules: depth estimation of camera image(37 ms) and LiDAR projection with depth completion (12 ms). As these operate in parallel, the latency of this stage is bottlenecked by the depth estimation process. Following this, the network inference and the PnP solver consume 29 ms and 25 ms, respectively. With an input resolution of 960x320, the total pipeline latency is approximately 91 ms, enabling a calibration throughput of about 11 Hz. This demonstrates that UniCalib achieves real-time performance, making it suitable for applications in practical scenarios such as autonomous driving.
5 Conclusion
In this paper, we presented UniCalib, a novel method that reformulates the cross-modal calibration task as a probabilistic intra-modality depth flow estimation problem. UniCalib bridges the modality gap by generating dense depth maps from both camera and LiDAR inputs, enabling a unified encoder to learn consistent features within a shared representation space. To address real-world uncertainty, our framework predicts a probabilistic depth flow field, guided by a reliability map and enhanced by the PWSF loss that adaptively suppresses supervision in unreliable regions. Experimental results on multiple datasets demonstrate that UniCalib achieves accurate extrinsic calibration with strong generalization capability. In future work, we plan to extend this probabilistic flow estimation framework to other cross-modal alignment problems, such as radar-camera calibration, where handling sparse and noisy measurements remains a key challenge.